# Detecting eclipsing double white dwarfs with electromagnetic and gravitational waves

### v1

@@ 5 Joint observation
In the previous two sections, we examined separately the capabilities of Gaia and LSST in optical observations and the capabilities of TianQin and LISA in GW detections for the DWDs in our Galaxy. In this section, we examine the properties of the DWDs detectable by both GW and EM observations.
In Table 7, we show the numbers of the DWDs with SNRs greater than 7 in Model 1 and Model 2 during the 5-year and 4-year nominal mission durations for TianQin and LISA, respectively. In addition, we count the numbers of eclipsing DWDs, constrained by the criteria such as sky area, apparent magnitude, and chi-square test, within Model 1 and Model 2 during the 5-year and 10-year nominal mission lifetimes of Gaia and LSST, respectively. The numbers of DWDs detectable by either EM or GW observations in Model 2 are about 2-3 times larger than in Model 1. The detection of DWDs by GW is influenced by a number of factors, including the source distance, spatial position, mass, orbital period, and so forth. The main factor contributing to the differences in the numbers of GW detections between Model 1 and Model 2 is the orbital period distribution of simulated DWDs. For example, in Model 2, the thin disk of our Galaxy follows a constant star formation history over a period of 10 billion years, resulting in more DWDs with frequencies in the millihertz range. On the other hand, we find that the main factor causing the difference in the numbers of EM detections between Model 1 and Model 2 is the effective surface temperature of the companion stars. We can see that the joint observation of TianQin and Gaia identifies 2 (Model 1) and 9 (Model 2) eclipsing DWDs. Similarly, the joint observation of TianQin and LSST identifies 7 (Model 1) and 11 (Model 2) eclipsing DWDs. In addition, the joint observation of LISA and Gaia identifies 4 (Model 1) and 29 (Model 2) eclipsing DWDs, while the joint observation of LISA and LSST identifies 15 (Model 1) and 47 (Model 2) eclipsing DWDs.
@@
Table 7: Number of DWDs from Model 1 (Model 2) that are detectable by TianQin, LISA, Gaia, LSST individually (diagonal elements), and their joint observation (off-diagonal elements).
TianQin LISA Gaia LSST
@@ 6 Summary and Conclusions
In this work, we simulated two populations of DWDs in our Galaxy with different star formation histories (Model 1 and Model 2) using binary population synthesis method. We predicted the numbers of DWDs detectable by TianQin, LISA, Gaia, and LSST individually and jointly in these two samples. We found that TianQin can detect 4969 (10180) DWDs with SNRs greater than 7 within a 5-year nominal mission duration in Model 1 (Model 2), while LISA can detect 16960 (32815) DWDs with SNRs greater than 7 within a 4-year nominal mission duration in Model 1 (Model 2). The number of DWDs detectable by TianQin and LISA in Model 2 is found to be twice that of in Model 1. The joint observation of TianQin and LISA can detect an additional 2,000 DWDs compared to LISA alone in Model 2. For EM observations, we found that Gaia and LSST can detect 67 (174) and 273 (554) compact eclipsing DWDs in Model 1 (Model 2) with orbital period less than 30 hours, respectively. The increased number in Model 2 is mainly due to the higher surface temperatures of the WDs generated in Model 2, making them more easily detectable by optical telescopes.
In the investigation of EM observations of eclipsing DWDs, we found that the $r$-band of LSST has superior performance compared to other bands, with a limiting magnitude of 24.7. The DWDs detectable by optical telescopes are predominantly distributed near the Galactic disk. Given that LSST primarily monitors the southern hemisphere, the inclusion of ZTF and other facilities in the northern hemisphere is crucial to compensate for the lack of sky coverage. In the near future, the Chinese Space Station Telescope (CSST) (Gong et al., 2019) with a $g$-band magnitude of 26.4 and a field of view of about 1.1 deg2 will have the potential to conduct deeper follow-up observations and find more DWDs reported by the space-borne GW detectors. Relevant work is currently under our investigation.
@@
For the space-borne detectors, we calculated the SNRs of DWDs for TianQin and LISA, individually and jointly. The latter can detect about $7\%$ more DWDs compared to the LISA alone. Furthermore, we analyzed the estimation errors of the parameters using the Fisher information matrix. From the cumulative distributions of the relative errors in Figure 10, we found that the main contributions to $\Delta{\mathcal{M}}/\mathcal{M}$ and $\Delta{\mathcal{D}}/\mathcal{D}$ are associated with $\Delta{\dot{f}}/\dot{f}$ and $\Delta\mathcal{A}/\mathcal{A}$. Due to the sensitive frequencies of the GW detectors, almost all DWDs detectable by LISA and TianQin are in the range of $10^{-3}-10^{-2}$ Hz, corresponding to orbital periods of 3-30 mins. In contrast, the orbital periods of DWDs observed in EM observations are mainly centered around one hour.
@@
In this work, we have considered only detached DWDs in our simulated population and have produced their light curves. In the future, we will study semi-detached DWDs and simulate their EM photometric and spectroscopic signatures. Furthermore, we plan to conduct direct comparison with existing and forthcoming EM surveys of DWDs like from the Zwicky Transient Facility (ZTF; Ren et al., 2023), Gaia, the Wide Field Survey Telescope (WFST; Wang et al., 2023), and the Multi-channel Photometric Survey Telescope (Mephisto; Yuan et al., 2020), to refine our population synthesis models.
@@ Acknowledgements
We thank Valeriya Korol, Shenghua Yu and Xiang-Dong Li for helpful discussions. HMJ is grateful for the hospitality of Nanjing University, where part of this work was done during his visit. YW gratefully acknowledges support from the National Key Research and Development Program of China (No. 2022YFC2205201, 2020YFC2201400), the National Natural Science Foundation of China (NSFC) under Grants No. 11973024, Major Science and Technology Program of Xinjiang Uygur Autonomous Region (No. 2022A03013-4), and Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001). YS acknowledges support from the National Key Research and Development Program of China (Grant Nos 2021YFA0718500 and 2023YFA1607902). BM acknowledges support from NSFC (12073092, 12103097). We acknowledge the High Performance Computing Platform at Huazhong University of Science and Technology for providing computational resources.
@@
Appendix A
{ThreePartTable}
Table 8: Physical properties of the known 16 eclipsing DWDs.
Source $R_{1}$ $R_{2}$ $m_{1}$ $m_{2}$ $\iota$ $P$ $d$ Refs.
($R_{\sun}$) ($R_{\sun}$) ($M_{\sun}$) ($M_{\sun}$) $(^{\circ})$ $(\mathrm{hour})$ $(\mathrm{pc})$
NLTT 11748 0.0423 ± 0.0004 0.0108 ± 0.0001 0.136 ± 0.007 0.707 ± 0.008 89.67 ± 0.12 5.64 150±32 1,2
CSS 41177 0.0210 ± 0.0026 0.0174 ± 0.0031 0.283 ± 0.064 0.274 ± 0.034 89.2 ± 0.3 2.78 350 ± 13 3
GALEX J1717+6757 0.101 ± 0.014 0.009 0.185 ± 0.010 0.90 86.9 ± 0.4 5.91 195±7 4,5
SDSS J065133.338+284423.37 0.0371 ± 0.0012 0.0142 ± 0.001 0.247 ± 0.015 0.49 ± 0.02 $86.9^{+1.6}_{-1.0}$ 0.21 1000.0 ± 100 6,7
SDSS J075141.18+014120.9 0.155 ± 0.020 0.0092 ± 0.0026 0.19 0.97+0.06 $85.4^{+4.2}_{-9.4}$ 1.90 1700±100 8
SDSS J115219.99+024814.4 $0.0223^{+0.0064}_{-0.0050}$ 0.0197 ± 0.0035 0.442 ± 0.088 0.47 ± 0.11 $89.3^{+0.140}_{-0.081}$ 2.40 $464^{+90}_{-85}$ 9
SDSS J082239.54+304857.19 $0.022^{+0.015}_{-0.014}$ $0.0041^{+0.0082}_{-0.0037}$ 0.304 ± 0.014 0.524 ± 0.050 $88.1^{+1.4}_{-2.3}$ 0.67 1110±80 10
ZTF J153932.16 + 502738.8 $0.01562^{+0.00038}_{-0.00038}$ $0.03140^{+0.00054}_{-0.00052}$ $0.610^{+0.017}_{-0.022}$ $0.210^{+0.014}_{-0.015}$ $84.15^{+0.64}_{-0.57}$ 0.12 2340 ± 140 7
ZTF J190125.42+530929.5 $0.029^{+0.001}_{-0.002}$ $0.022^{+0.003}_{-0.002}$ $0.36^{+0.04}_{-0.04}$ $0.36^{+0.05}_{-0.05}$ $86.2^{+0.6}_{-0.2}$ 0.68 $915^{+96}_{-80}$ 11,12
ZTF J053802.73+195302.89 $0.02069_{-0.00034}^{+0.0028}$ $0.02319_{-0.00026}^{+0.00032}$ $0.45_{-0.05}^{+0.05}$ $0.32_{-0.03}^{+0.03}$ $85.43^{+0.07}_{0.09}$ 0.2407 $684^{+18}_{-14}$ 12
ZTF J202922.31+153430.97 $0.029_{-0.003}^{+0.002}$ $0.028_{-0.003}^{+0.003}$ $0.32_{-0.04}^{+0.04}$ $0.30_{-0.04}^{+0.04}$ $86.64^{0.70}_{0.4}$ 0.3478 $2020^{+150}_{-150}$ 12
ZTF J072221.49+183930.57 $0.0224_{-0.0002}^{+0.0004}$ $0.0249_{-0.0003}^{+0.0001}$ $0.38_{-0.04}^{+0.04}$ $0.33_{-0.03}^{+0.03}$ $89.66^{+0.22}_{-0.22}$ 0.3950 $928^{+23}_{-31}$ 12
ZTF J174955.30+092432.40 $0.022_{-0.004}^{+0.003}$ $0.025_{-0.004}^{+0.004}$ $0.40_{-0.05}^{+0.07}$ $0.28_{-0.04}^{+0.05}$ $85.45^{1.40}_{1.15}$ 0.4405 $1550^{+200}_{-180}$ 12
ZTF J2243+5242 $0.0298^{+0.0013}_{-0.0012}$ $0.0275^{+0.0012}_{-0.0013}$ $0.323^{+0.065}_{0.047}$ $0.335^{+0.052}_{0.054}$ $82.12^{+0.51}_{-0.38}$ 0.1466 $2425^{+108}_{-114}$ 13
J0225+6920 0.0291 ± 0.001 0.0244 ± 0.0007 0.4 ± 0.04 0.28 ± 0.02 85.25 ± 0.06 0.7865 $402.6^{+5.9}_{-6.8}$ 14
J2102+4145 0.0211 ± 0.0002 $0.0203^{+0.0002}_{0.0003}$ 0.375 ± 0.003 0.314 ± 0.003 $88.693^{+0.006}_{-0.005}$ 2.4048 165 ± 1 15,16
  • [1] Kaplan et al. (2014). [2] Steinfadt et al. (2010). [3] Parsons et al. (2011). [4] Vennes et al. (2011). [5] Hermes et al. (2014). [6] Brown et al. (2011). [7] Burdge et al. (2019). [8] Kilic et al. (2014). [9] Hallakoun et al. (2016). [10] Brown et al. (2017). [11] Coughlin et al. (2020). [12] Burdge et al. (2020a). [13] Burdge et al. (2020b). [14] Munday et al. (2023) [15] Kosakowski et al. (2023) [16] Amaral et al. (2024)

### v2

@@ 5 Joint observation
In the previous two sections, we examined separately the capabilities of Gaia and VRO in optical observations and the capabilities of TianQin and LISA in GW detections for the DWDs in our Galaxy. In this section, we examine the properties of the DWDs detectable by both GW and EM observations.
In Table 7, we show the numbers of the DWDs with SNRs greater than 7 in Model 1 and Model 2 during the 5-year and 4-year nominal mission durations for TianQin and LISA, respectively. In addition, we count the numbers of eclipsing DWDs, constrained by the criteria such as sky area, apparent magnitude, and chi-square test, within Model 1 and Model 2 during the 5-year and 10-year nominal mission lifetimes of Gaia and VRO, respectively. The numbers of DWDs detectable by either EM or GW observations in Model 2 are about 2-3 times larger than in Model 1. The detection of DWDs by GW is influenced by a number of factors, including the source distance, spatial position, mass, orbital period, and so forth. The main factor contributing to the differences in the numbers of GW detections between Model 1 and Model 2 is the orbital period distribution of simulated DWDs. For example, in Model 2, the thin disk of our Galaxy follows a constant star formation history over a period of 10 billion years, resulting in more DWDs with frequencies in the millihertz range. On the other hand, we find that the main factor causing the difference in the numbers of EM detections between Model 1 and Model 2 is the effective surface temperature of the companion stars. We can see that the joint observation of TianQin and Gaia identifies 2 (Model 1) and 9 (Model 2) eclipsing DWDs. Similarly, the joint observation of TianQin and VRO identifies 7 (Model 1) and 11 (Model 2) eclipsing DWDs. In addition, the joint observation of LISA and Gaia identifies 4 (Model 1) and 29 (Model 2) eclipsing DWDs, while the joint observation of LISA and VRO identifies 15 (Model 1) and 47 (Model 2) eclipsing DWDs.
@@
Table 7: Number of DWDs from Model 1 (Model 2) that are detectable by TianQin, LISA, Gaia, VRO individually (diagonal elements), and their joint observation (off-diagonal elements).
TianQin LISA Gaia VRO
@@ 6 Summary and Conclusions
In this work, we simulated two populations of DWDs in our Galaxy with different star formation histories (Model 1 and Model 2) using the binary population synthesis method. Model 1 assumes a burst of star formation 13.7 Gyr ago, while Model 2 assumes a constant star formation rate for the thin disk and a burst of star formation for the thick disk and bulge. We predicted the numbers of DWDs detectable by TianQin, LISA, Gaia, and VRO individually and jointly in these two samples. We found that TianQin can detect 4969 (10180) DWDs with SNRs greater than 7 within a 5-year nominal mission duration in Model 1 (Model 2), while LISA can detect 16960 (32815) DWDs with SNRs greater than 7 within a 4-year nominal mission duration in Model 1 (Model 2). The number of DWDs detectable by TianQin and LISA in Model 2 is found to be twice that of in Model 1. The joint observation of TianQin and LISA can detect an additional 2,000 DWDs compared to LISA alone in Model 2. For EM observations, we found that Gaia and VRO can detect 67 (174) and 273 (554) compact eclipsing DWDs in Model 1 (Model 2) with orbital period less than 30 hours, respectively. The increased number in Model 2 is mainly due to the higher surface temperatures of the WDs generated in Model 2, making them more easily detectable by optical telescopes.
@@
For the space-borne detectors, we calculated the SNRs of DWDs for TianQin and LISA, individually and jointly. The latter can detect about $7\%$ more DWDs compared to the LISA alone. Furthermore, we analyzed the estimation errors of the parameters using the Fisher information matrix. Due to the sensitive frequencies of the GW detectors, almost all DWDs detectable by LISA and TianQin are in the range of $10^{-3}-10^{-2}$ Hz, corresponding to orbital periods of 3-30 mins. In contrast, the orbital periods of DWDs observed in EM observations are mainly centered around one hour.
In the investigation of EM observations of eclipsing DWDs, we found that the $r$-band of VRO has superior performance compared to other bands, with a limiting magnitude of 24.7. The DWDs detectable by optical telescopes are predominantly distributed near the Galactic disk.
In this work, we have focused solely on simulated observations from VRO and Gaia, but it is important to recognize that many other survey projects will also contribute to the detection of more DWDs. Since VRO primarily observes the southern hemisphere, including surveys such as ZTF and other northern hemisphere facilities is essential to achieve broader sky coverage. For example, the Chinese Space Station Telescope (CSST) (Gong et al., 2019) with a $g$-band magnitude of 26.4 and a field of view of about 1.1 deg2 is expected to identify more DWDs candidates during its primary survey mission. The current sample of known DWDs is obtained through data mining across a combination of photometric, spectral, and astrometric sky surveys, supplemented by dedicated follow-up observations. To detect more DWDs using EM methods in the future, a similar approach will be necessary—even in the VRO era—by first identifying DWD candidates from large-scale surveys and then conducting targeted follow-up observations using both ground-based and space-based telescopes.
@@
Furthermore, much works remain to be done in the near future. For instance, our current simulation only considers the detached DWDs, and we plan to extend our study to include semi-detached DWDs. Additionally, direct comparisons between our population synthesis results and both existing and forthcoming EM surveys of DWDs, such as those from the Zwicky Transient Facility (ZTF; Ren et al., 2023), Gaia, the Wide Field Survey Telescope (WFST; Wang et al., 2023), and the Multi-channel Photometric Survey Telescope (Mephisto; Yuan et al., 2020) will be crucial in refining our binary population synthesis models and improving their accuracy.
@@ Acknowledgements
We thank Valeriya Korol, Shenghua Yu and Xiang-Dong Li for helpful discussions. HMJ is grateful for the hospitality of Nanjing University, where part of this work was done during his visit. YW gratefully acknowledges support from the National Key Research and Development Program of China (No. 2023YFC2206702 and No. 2022YFC2205201), the National Natural Science Foundation of China (NSFC) under Grants No. 11973024, Major Science and Technology Program of Xinjiang Uygur Autonomous Region (No. 2022A03013-4), and Guangdong Major Project of Basic and Applied Basic Research (No. 2019B030302001). YS acknowledges support from the National Key Research and Development Program of China (No. 2021YFA0718500 and No. 2023YFA1607902). BM acknowledges support from NSFC (No. 12073092 and No. 12103097). We acknowledge the High Performance Computing Platform at Huazhong University of Science and Technology for providing computational resources. The authors thank the anonymous referee for helpful comments and suggestions that significantly improved our manuscript.